# Learning Interactive Real-World Simulators

### v2

@@ 4 Applications of the UniSim
Language models and vision language models (VLM) have recently been used as policies that can operate in image or text based observation and action spaces (du2023guiding, ; driess2023palm, ; brohan2023rt, ). One major challenge in learning such agents lies in the need for large amounts of language action labels. The labor intensity in data collection only increases as tasks increase in horizon and complexity. The UniSim can generate large amounts of training data for VLM policies through hindsight relabeling.
Setup and Baseline. We use data from the Language Table environment (lynch2020language, ) for learning geometric rearrangements of blocks on a table. We train an image-goal conditioned VLM policy to predict language instructions and the motor controls from the start and goal images using the PALM-E architecture (driess2023palm, ) (See data and model details in Appendix 11.1). For the baseline, the goal is set to the last frame of the original short-horizon trajectories. During each evaluation run, we set the long-horizon goal by modifying the location of 3-4 blocks, and measure the blocks’ distance to their goal states after executing 5 instructions using the VLM policy. We define the reduction in distance to goal (RDG) metric as
@@
Generating Hindsight Data with the Simulator. To use the simulator for long-horizon tasks, we draw inspiration from hindsight relabeling (rauber2017hindsight, ). Specifically, we create a total of 10k long-horizon trajectories from the simulator by doing rollouts in the simulator 3-5 times per trajectory, where each rollout corresponds to one scripted language instruction. We then use the final frame from each long-horizon rollout as a goal input and the scripted language instructions as supervision for training the VLM policy.
Results on Real-Robot Evaluation. Despite the VLM policy only being trained on simulated data, it is able to produce effective high-level language actions given an initial and goal image from the real Language Table domain where the data for training the simulator was collected. The simulator can simulate video trajectories from the initial real observation, from which robot actions are recovered using an inverse dynamics model and executed on the real robot. Figure 7 shows that the language actions produced by the VLM, the generated videos from the simulator according to the language actions, and the executions on the real robot. We see that the simulated video trajectory is successfully translated to robot actions in the real world. See additional results from the long-horizon VLM policy in Appendix 8.2.
@@ 4.2 Real-World Simulator for Reinforcement Learning
Reinforcement learning (RL) has achieved superhuman performance on difficult tasks such as playing Go and Atari games (silver2017alphazero, ; mnih2015human, ), but has limited real world applications due, among other reasons, to the lack of a realistic environment simulator (dulac2019challenges, ). We investigate whether the simulator can enable effective training of RL agents by providing the agent with a realistic simulator that can be accessed in parallel.
Setup. We finetune the PaLI 3B vision-language model (chen2022pali, ) to predict low-level control actions (joint movements in $\Delta x,\Delta y$) from an image observation and a task description (e.g., “move the blue cube to the right”) using behavioral cloning (BC) to serve as the low-level control policy and the baseline, which we call the vision-language-action (VLA) policy similar to brohan2023rt . Because the UniSim can take low-level control actions as input, we can directly conduct model-based rollouts in the simulator using control actions generated by VLA policy. To acquire reward information, we use the number of steps-to-completion from the training data as a proxy reward to train a model that maps the current observation to learned reward. We then use the REINFORCE algorithm (williams1992simple, ) to optimize the VLA policy, treating the rollouts from the simulator as the on-policy rollouts from the real environment and use the learned reward model to predict rewards from simulated rollouts. See details of RL training in Appendix 11.2.
Figure 8: [Top] Simulation from low-level controls. The UniSim supports low-level control actions as inputs to move endpoint horizontally, vertically, and diagonally. [Bottom] Real-robot execution of an RL policy trained in simulation and zero-shot onto the real Language Table task. The RL policy can successfully complete the task of “moving blue cube to green circle”.
Results. We first do a sanity check on simulating real-robot executions by applying low-level control actions (e.g., $\Delta x=0.05,\delta y=0.05$) repeatedly for 20-30 environment steps to move the endpoint left, right, down, up, and diagonally in Figure 8 (top two rows). We see that the simulated rollouts capture both the endpoint movements and the physics of collision. To compare the RL policy trained in simulation to the BC policy, we qualitatively assessed the simulated rollouts in the simulator. Table 3 shows that RL training significantly improves the performance of the VLA policy across a wide set of tasks, especially in tasks such as “point to blue block”. We then directly deploy the RL policy trained in the simulator onto the real robot in zero-shot, and observe successful task executions as shown in Figure 8 (bottom row). Additional results on real robot can be found in Appendix 8.3.
@@ 4.3 Realistic Simulator for Broader Vision-Language Tasks
Table 4: VLM trained in the UniSim to perform video captioning tasks. CIDEr scores for PaLI-X finetuned only on simulated data from the UniSim compared to no finetuning and finetuning on true video data from ActivityNet Captions. Finetuning only on simulated data has a large advantage over no finetuning and transfers better to other tasks than finetuning on true data.
The UniSim can generate training data for other machine-learning subproblems. This is especially useful when natural data is rare or difficult to collect (e.g., footage of crimes or accidents). We provide such a proof-of-concept by training vision-language models on purely generated data from the UniSim, and observe significant performance benefits in video captioning.
Setup. We finetune PaLI-X (chen2023pali, ), a VLM with 55B parameters pretrained on a broad set of image, video, and language tasks, to caption a set of videos generated by the UniSim using texts from the training split of ActivityNet Captions (krishna2017dense, ). We measure the CIDEr score of the finetuned model on the test split of ActivityNet Captions as well as other captioning tasks following the same setup as chen2023pali . See finetuning details of PaLI-X in Appendix 11.3.
Results. We compare PaLI-X finetuned on purely generated videos to pretrained PaLI-X without finetuning and PaLI-X finetuned on original ActivityNet Captions in Table 4. Purely finetuning on generated data drastically improves the captioning performance from no finetuning at all on ActivityNet (15.2 to 46.23), while achieving 84% performance of finetuning on true data. Furthermore, PaLI-X finetuned on generated data transfers better to other captioning tasks such as MSR-VTT (xu2016msr, ), VATEX (wang2019vatex, ), and SMIT (monfort2021spoken, ) than PaLI-X finetuned on true data, which tends to overfit to ActivityNet. These results suggest that the UniSim can serve as an effective data generator for improving broader vision-language models.
@@ 5 Related Work
Internet-Scale Generative Models. Language models trained on internet text succeed at text-based tasks (openai2023gpt4, ; anil2023palm, ) but not physical tasks, which requires perception and control. Internet-scale generative models can synthesize realistic images and videos (wu2021godiva, ; ho2022imagen, ; singer2022make, ; yang2023probabilistic, ; blattmann2023align, ), but have mostly been applied to generative media (zhang2023text, ) as opposed to empowering sophisticated agents capable of multi-turn interactions. du2302learning shows video generation can serve as policies, but the major bottleneck for policy learning often lies in limited access to real-world environments (dulac2019challenges, ). We focus on this exact bottleneck by learning universal simulators of the real world, enabling realistic and unlimited “environment” access for training sophisticated agents interactively.
Learning World Models. Learning an accurate dynamics model in reaction to control inputs has been a long-standing challenge in system identification (ljung1994modeling, ), model-based reinforcement learning (sutton1991dyna, ), and optimal control AstromWittenmark1973 ; bertsekas1995dynamic . Most systems choose to learn one dynamics model per system in the lower dimensional state space as opposed to in the pixel space (ferns2004metrics, ; achille2018separation, ; lesort2018state, ; castro2020scalable, ), which, despite being a simpler modeling problem, limits knowledge sharing across systems. With large transformer architectures, learning image-based world models has become plausible (hafner2020mastering, ; chen2022transdreamer, ; seo2022reinforcement, ; micheli2022transformers, ; wu2022slotformer, ; hafner2023mastering, ), but mostly in games or simulated domains with visually simplistic and abundant data. In generative modeling of videos, previous works have leveraged text prompts (yu2023video, ; zhou2022magicvideo, ), driving motions (siarohin2019animating, ; wang2022latent, ), 3D geometries (weng2019photo, ; xue2018visual, ), physical simulations (chuang2005animating, ), frequency information (li2023generative, ), and user annotations (hao2018controllable, ) to introduce movements into videos. However, they focus on generating domain specific videos (e.g., for self-driving) as opposed to building a universal simulator that can be used to further improve other agents. The amount of control over generated videos in these existing work is also limited, as they do not treat video generation as a dynamics modeling problem like in our work.
@@ 6 Limitations and Conclusion
We have shown it is possible to learn a simulator of the real world in response to various action inputs ranging from texts to robot controls. The UniSim can simulate visually realistic experiences for interacting with humans and training autonomous agents. We hope the UniSim will instigate broad interest in learning and applying real-world simulators to improve machine intelligence. Our simulator has a few limitations that call for future work:
@@
7 Acknowledgements
We thank Hanjun Dai and Doina Precup for reviewing this manuscript. We gratefully acknowledges the support of a Canada CIFAR AI Chair, NSERC and Amii, NSF GRFP, and the support from Berkeley BAIR industrial consortion.
@@ 7 Acknowledgements
In this Appendix we provide additional qualitative results on long-horizon simulation of human and robot interactions (Section 8.1), long-horizon VLM policies (Section 8.2), and low-level RL policies (Section 8.3) that work on real robot. We also provided details on the dataset used to train UniSim in Section 9, the model architecture and training details of UniSim in Section 10, and the details of the three experimental setups for applications of UniSim in Section 11. Finally, we provide failed examples when UniSim is not jointly trained on broad datasets (Section 13). Video demos can be found at anonymous-papers-submissions.github.io
8 Additional Results
8.1 Additional Long-Horizon Interaction
@@
8.2 Additional Real-Robot Results for Long-Horizon Language Policy
@@
8.3 Additional Results on Learning RL Policy in UniSim
@@
9 Datasets
We provide the datasets used to train UniSim below, including dataset name, number of training examples (approximate), and weight in the data mixture. Miscellaneous data are collections of datasets that have not been published. Some of these datasets have been processed into train and validation split, hence the number of training examples may differ from the original data size. When text are available in the original dataset, we use T5 language model embeddings [24] to preprocess the text into continuous representations. When low-level controls are available in the original dataset, we encode them both as text and normalize then discretize them into 4096 bins contatenated with language embeddings (if present). The choice of mixture weights are either 0.1 or 0.05 without careful tuning. How data mixture weights affect simulation performance is an interesting line of future work.
@@ 9 Datasets
Simulation Habitat HM3D [16] 710 0.1
Language Table sim [33] 160k 0.05
Real Robot Bridge Data [18] 2k 0.05
RT-1 data [19] 70k 0.1
Language Table real [33] 440k 0.05
@@
Human activities Ego4D [9] 3.5M 0.1
Something-Something V2 [21] 160k 0.1
EPIC-KITCHENS [20] 25k 0.1
@@
Panorama scan Matterport Room-to-Room scans [73] 3.5M 0.1
Internet text-image LAION-400M [23] 400M 0.05
ALIGN [74] 400M 0.05
@@
10 Architecture and Training
We the 3D U-Net architecture [75, 28] to parametrize the UniSim video model. We apply the spatial downsampling pass followed by the spatial upsampling pass with skip connections to the downsampling pass activations with interleaved 3D convolution and attention layers as in the standard 3D U-Net. The video models in UniSim consist of one history conditioned video prediction model as the base and two additional spatial super-resolution models similar to  ho2022imagen [3]. The history conditioned base model operates at temporal and spatial resolution $[16,24,40]$, and the two spatial super-resolution models operate at spatial resolution $[24,40]\rightarrow[48,80]$ and $[48,80]\rightarrow[192,320]$, respectively. To condition the base video model on the history, we take 4 frames from the previous video segment and concatenate them channelwise to the noise samples inputted to the U-Net. We employ temporal attention for the forward model to allow maximum modeling flexibility but temporal convolution to the super-resolution models for efficiency reasons similar to ho2022imagen [3]. The model and training hyperparamters of UniSim are summarized in Table 6.
@@ 10 Architecture and Training
Table 6: Hyperparameters for training the UniSim diffusion model.
11 Details of Experimental Setups
11.1 Details of Learning Long-Horizon Policy
@@
Language Table Dataset and environment. The Language Table [33] dataset consists of 160k simulated trajectories and 440k real trajectories where each trajectory contains a language instruction (e.g., “move blue cube to the right”), a sequence of visuomotor controls, and a sequence of image frames corresponding to the execution of the task. The original trajectories have short horizons (e.g., only moving one block).
@@
PALM-E VLM Policy. We modify the original PALM-E 12B model [31] to condition on a goal image as additional input before decoding the text actions. The VLM is finetuned on either the original short horizon data or the long horizon simulated data using 64 TPUv3 chips for 1 day. The supervision for short-horizon baseline is the single step language instruction in the original data, whereas the supervision for long-horizon UniSim data is the scripted long-horizon language instructions chained together that generated the video data. Other model architecture and training details follow driess2023palm [31].
@@
Simulated evaluation. In setting up goal in the simulated environments, a subset of 3-4 blocks (randomly selected) are moved by 0.05, 0.1, or 0.2 along the x,y axes (randomly selected). The original observation space has $x\in[0.15,0.6]$ and $y\in[-0.3048,0.3048]$. So the modification of goal location corresponds to meaningful block movements. For executing the long-horizon VLM policy trained on UniSim data, we first sample one language instruction from the VLM, predict a video of 16 frames, and use a separately trained inverse dynamics model similar to  du2302learning [49] to recover the low-level control actions, which we found to slightly outperform directly regressing on control actions from language outputs of the VLM. We execute 5 instructions in total, and measure the final distance to goal according to the ground truth simulator state. We 5 evaluations each with a different random seed for sampling the initial state and resetting the goal, and report the mean and standard error in Table 3.
11.2 Details of RL Policy Training

### v3

@@ 4 Applications of UniSim
Language models and vision language models (VLM) have recently been used as policies that can operate in image or text based observation and action spaces (Du et al., 2023b; Driess et al., 2023; Brohan et al., 2023). One major challenge in learning such agents lies in the need for large amounts of language action labels. The labor intensity in data collection only increases as tasks increase in horizon and complexity. UniSim can generate large amounts of training data for VLM policies through hindsight relabeling.
Setup and Baseline. We use data from the Language Table environment (Lynch & Sermanet, 2020) for learning geometric rearrangements of blocks on a table. We train an image-goal conditioned VLM policy to predict language instructions and the motor controls from the start and goal images using the PALM-E architecture (Driess et al., 2023) (See data and model details in Appendix D.1). For the baseline, the goal is set to the last frame of the original short-horizon trajectories. During each evaluation run, we set the long-horizon goal by modifying the location of 3-4 blocks, and measure the blocks’ distance to their goal states after executing 5 instructions using the VLM policy. We define the reduction in distance to goal (RDG) metric as
@@
Generating Hindsight Data with the Simulator. To use the simulator for long-horizon tasks, we draw inspiration from hindsight relabeling (Rauber et al., 2019). Specifically, we create a total of 10k long-horizon trajectories from the simulator by doing rollouts in the simulator 3-5 times per trajectory, where each rollout corresponds to one scripted language instruction. We then use the final frame from each long-horizon rollout as a goal input and the scripted language instructions as supervision for training the VLM policy.
Results on Real-Robot Evaluation. Despite the VLM policy only being trained on simulated data, it is able to produce effective high-level language actions given an initial and goal image from the real Language Table domain where the data for training the simulator was collected. The simulator can simulate video trajectories from the initial real observation, from which robot actions are recovered using an inverse dynamics model and executed on the real robot. Figure 7 shows that the language actions produced by the VLM, the generated videos from the simulator according to the language actions, and the executions on the real robot. We see that the simulated video trajectory is successfully translated to robot actions in the real world. See additional results from the long-horizon VLM policy in Appendix A.2.
@@ 4.2 Real-World Simulator for Reinforcement Learning
Reinforcement learning (RL) has achieved superhuman performance on difficult tasks such as playing Go and Atari games (Silver et al., ; Mnih et al., 2015), but has limited real world applications due, among other reasons, to the lack of a realistic environment simulator (Dulac-Arnold et al., 2019). We investigate whether the simulator can enable effective training of RL agents by providing the agent with a realistic simulator that can be accessed in parallel.
Setup. We finetune the PaLI 3B vision-language model (Chen et al., 2022b) to predict low-level control actions (joint movements in $\Delta x,\Delta y$) from an image observation and a task description (e.g., “move the blue cube to the right”) using behavioral cloning (BC) to serve as the low-level control policy and the baseline, which we call the vision-language-action (VLA) policy similar to Brohan et al. (2023). Because UniSim can take low-level control actions as input, we can directly conduct model-based rollouts in the simulator using control actions generated by VLA policy. To acquire reward information, we use the number of steps-to-completion from the training data as a proxy reward to train a model that maps the current observation to learned reward. We then use the REINFORCE algorithm (Williams, 1992) to optimize the VLA policy, treating the rollouts from the simulator as the on-policy rollouts from the real environment and use the learned reward model to predict rewards from simulated rollouts. See details of RL training in Appendix D.2.
Figure 8: [Top] Simulation from low-level controls. UniSim supports low-level control actions as inputs to move endpoint horizontally, vertically, and diagonally. [Bottom] Real-robot execution of an RL policy trained in simulation and zero-shot onto the real Language Table task. The RL policy can successfully complete the task of “moving blue cube to green circle”.
Results. We first do a sanity check on simulating real-robot executions by applying low-level control actions (e.g., $\Delta x=0.05,\delta y=0.05$) repeatedly for 20-30 environment steps to move the endpoint left, right, down, up, and diagonally in Figure 8 (top two rows). We see that the simulated rollouts capture both the endpoint movements and the physics of collision. To compare the RL policy trained in simulation to the BC policy, we qualitatively assessed the simulated rollouts in the simulator. Table 3 shows that RL training significantly improves the performance of the VLA policy across a wide set of tasks, especially in tasks such as “point to blue block”. We then directly deploy the RL policy trained in the simulator onto the real robot in zero-shot, and observe successful task executions as shown in Figure 8 (bottom row). Additional results on real robot can be found in Appendix A.3.
@@ 4.3 Realistic Simulator for Broader Vision-Language Tasks
Table 4: VLM trained in UniSim to perform video captioning tasks. CIDEr scores for PaLI-X finetuned only on simulated data from UniSim compared to no finetuning and finetuning on true video data from ActivityNet Captions. Finetuning only on simulated data has a large advantage over no finetuning and transfers better to other tasks than finetuning on true data.
UniSim can generate training data for other machine-learning subproblems. This is especially useful when natural data is rare or difficult to collect (e.g., footage of crimes or accidents). We provide such a proof-of-concept by training vision-language models on purely generated data from UniSim, and observe significant performance benefits in video captioning.
Setup. We finetune PaLI-X (Chen et al., 2023), a VLM with 55B parameters pretrained on a broad set of image, video, and language tasks, to caption a set of videos generated by UniSim using texts from the training split of ActivityNet Captions (Krishna et al., 2017). We measure the CIDEr score of the finetuned model on the test split of ActivityNet Captions as well as other captioning tasks following the same setup as Chen et al. (2023). See finetuning details of PaLI-X in Appendix D.3.
Results. We compare PaLI-X finetuned on purely generated videos to pretrained PaLI-X without finetuning and PaLI-X finetuned on original ActivityNet Captions in Table 4. Purely finetuning on generated data drastically improves the captioning performance from no finetuning at all on ActivityNet (15.2 to 46.23), while achieving 84% performance of finetuning on true data. Furthermore, PaLI-X finetuned on generated data transfers better to other captioning tasks such as MSR-VTT (Xu et al., 2016), VATEX (Wang et al., 2019), and SMIT (Monfort et al., 2021) than PaLI-X finetuned on true data, which tends to overfit to ActivityNet. These results suggest that UniSim can serve as an effective data generator for improving broader vision-language models.
@@ 5 Related Work
Internet-Scale Generative Models. Language models trained on internet text succeed at text-based tasks (OpenAI, 2023; Anil et al., 2023) but not physical tasks, which requires perception and control. Internet-scale generative models can synthesize realistic images and videos (Wu et al., 2021; Ho et al., 2022a; Singer et al., 2022; Yang et al., 2023; Blattmann et al., 2023), but have mostly been applied to generative media (Zhang et al., 2023) as opposed to empowering sophisticated agents capable of multi-turn interactions. Du et al. (2023a) shows video generation can serve as policies, but the major bottleneck for policy learning often lies in limited access to real-world environments (Dulac-Arnold et al., 2019). We focus on this exact bottleneck by learning universal simulators of the real world, enabling realistic and unlimited “environment” access for training sophisticated agents interactively.
Learning World Models. Learning an accurate dynamics model in reaction to control inputs has been a long-standing challenge in system identification (Ljung & Glad, 1994), model-based reinforcement learning (Sutton, 1991), and optimal control Åström & Wittenmark (1973); Bertsekas (1995). Most systems choose to learn one dynamics model per system in the lower dimensional state space as opposed to in the pixel space (Ferns et al., 2004; Achille & Soatto, 2018; Lesort et al., 2018; Castro, 2020), which, despite being a simpler modeling problem, limits knowledge sharing across systems. With large transformer architectures, learning image-based world models has become plausible (Hafner et al., 2020; Chen et al., 2022a; Seo et al., 2022; Micheli et al., 2022; Wu et al., 2022; Hafner et al., 2023), but mostly in games or simulated domains with visually simplistic and abundant data. In generative modeling of videos, previous works have leveraged text prompts (Yu et al., 2023; Zhou et al., 2022), driving motions (Siarohin et al., 2019; Wang et al., 2022), 3D geometries (Weng et al., 2019; Xue et al., 2018), physical simulations (Chuang et al., 2005), frequency information (Li et al., 2023), and user annotations (Hao et al., 2018) to introduce movements into videos. However, they focus on generating domain specific videos (e.g., for self-driving) as opposed to building a universal simulator that can be used to further improve other agents. The amount of control over generated videos in these existing work is also limited, as they do not treat video generation as a dynamics modeling problem like in our work.
@@ 6 Limitations and Conclusion
We have shown it is possible to learn a simulator of the real world in response to various action inputs ranging from texts to robot controls. UniSim can simulate visually realistic experiences for interacting with humans and training autonomous agents. We hope UniSim will instigate broad interest in learning and applying real-world simulators to improve machine intelligence. Our simulator has a few limitations that call for future work:
@@
In this Appendix we provide additional qualitative results on long-horizon simulation of human and robot interactions (Section A.1), long-horizon VLM policies (Section A.2), and low-level RL policies (Section A.3) that work on real robot. We also provided details on the dataset used to train UniSim in Section B, the model architecture and training details of UniSim in Section C, and the details of the three experimental setups for applications of UniSim in Section D. Finally, we provide failed examples when UniSim is not jointly trained on broad datasets (Section F). Video demos can be found at anonymous-papers-submissions.github.io
Appendix A Additional Results
A.1 Additional Long-Horizon Interaction
@@
A.2 Additional Real-Robot Results for Long-Horizon Language Policy
@@
A.3 Additional Results on Learning RL Policy in UniSim
@@
Appendix B Datasets
We provide the datasets used to train UniSim below, including dataset name, number of training examples (approximate), and weight in the data mixture. Miscellaneous data are collections of datasets that have not been published. Some of these datasets have been processed into train and validation split, hence the number of training examples may differ from the original data size. When text are available in the original dataset, we use T5 language model embeddings (Raffel et al., 2020) to preprocess the text into continuous representations. When low-level controls are available in the original dataset, we encode them both as text and normalize then discretize them into 4096 bins contatenated with language embeddings (if present). The choice of mixture weights are either 0.1 or 0.05 without careful tuning. How data mixture weights affect simulation performance is an interesting line of future work.
@@ Appendix B Datasets
Simulation Habitat HM3D (Ramakrishnan et al., 2021) 710 0.1
Language Table sim (Lynch & Sermanet, 2020) 160k 0.05
Real Robot Bridge Data (Ebert et al., 2021) 2k 0.05
RT-1 data (Brohan et al., 2022) 70k 0.1
Language Table real (Lynch & Sermanet, 2020) 440k 0.05
@@
Human activities Ego4D (Grauman et al., 2022) 3.5M 0.1
Something-Something V2 (Goyal et al., 2017) 160k 0.1
EPIC-KITCHENS (Damen et al., 2018) 25k 0.1
@@
Panorama scan Matterport Room-to-Room scans (Anderson et al., 2018) 3.5M 0.1
Internet text-image LAION-400M (Schuhmann et al., 2021) 400M 0.05
ALIGN (Jia et al., 2021) 400M 0.05
@@
Appendix C Architecture and Training
We the 3D U-Net architecture (Çiçek et al., 2016; Ho et al., 2022b) to parametrize UniSim video model. We apply the spatial downsampling pass followed by the spatial upsampling pass with skip connections to the downsampling pass activations with interleaved 3D convolution and attention layers as in the standard 3D U-Net. The video models in UniSim consist of one history conditioned video prediction model as the base and two additional spatial super-resolution models similar to  Ho et al. (2022a). The history conditioned base model operates at temporal and spatial resolution $[16,24,40]$, and the two spatial super-resolution models operate at spatial resolution $[24,40]\rightarrow[48,80]$ and $[48,80]\rightarrow[192,320]$, respectively. To condition the base video model on the history, we take 4 frames from the previous video segment and concatenate them channelwise to the noise samples inputted to the U-Net. We employ temporal attention for the forward model to allow maximum modeling flexibility but temporal convolution to the super-resolution models for efficiency reasons similar to Ho et al. (2022a). The model and training hyperparamters of UniSim are summarized in Table 6.
@@ Appendix C Architecture and Training
Table 6: Hyperparameters for training UniSim diffusion model.
Appendix D Details of Experimental Setups
D.1 Details of Learning Long-Horizon Policy
@@
Language Table Dataset and environment. The Language Table (Lynch & Sermanet, 2020) dataset consists of 160k simulated trajectories and 440k real trajectories where each trajectory contains a language instruction (e.g., “move blue cube to the right”), a sequence of visuomotor controls, and a sequence of image frames corresponding to the execution of the task. The original trajectories have short horizons (e.g., only moving one block).
@@
PALM-E VLM Policy. We modify the original PALM-E 12B model (Driess et al., 2023) to condition on a goal image as additional input before decoding the text actions. The VLM is finetuned on either the original short horizon data or the long horizon simulated data using 64 TPUv3 chips for 1 day. The supervision for short-horizon baseline is the single step language instruction in the original data, whereas the supervision for long-horizon UniSim data is the scripted long-horizon language instructions chained together that generated the video data. Other model architecture and training details follow Driess et al. (2023).
@@
Simulated evaluation. In setting up goal in the simulated environments, a subset of 3-4 blocks (randomly selected) are moved by 0.05, 0.1, or 0.2 along the x,y axes (randomly selected). The original observation space has $x\in[0.15,0.6]$ and $y\in[-0.3048,0.3048]$. So the modification of goal location corresponds to meaningful block movements. For executing the long-horizon VLM policy trained on UniSim data, we first sample one language instruction from the VLM, predict a video of 16 frames, and use a separately trained inverse dynamics model similar to  Du et al. (2023a) to recover the low-level control actions, which we found to slightly outperform directly regressing on control actions from language outputs of the VLM. We execute 5 instructions in total, and measure the final distance to goal according to the ground truth simulator state. We 5 evaluations each with a different random seed for sampling the initial state and resetting the goal, and report the mean and standard error in Table 3.
D.2 Details of RL Policy Training